OpenStudy (anonymous):

suppose that (x+k)^2=x^2-2x+k^2 for all values of x. k=?

OpenStudy (anonymous):

k=-1 i think

OpenStudy (ash2326):

(x-1)^2=x^2-2x+(-1)^2 so k=-1